# IMPACT: A Generic Semantic Loss for Multimodal Medical Image Registration

### v2

@@ II-A Image Registration: An Optimization Problem
In unsupervised intensity-based image registration, the objective is to estimate a spatial transformation $T_{\theta}$: $\Omega_{F}\rightarrow\Omega_{M}$ that aligns a moving image $I_{M}$ : $\Omega_{M}\rightarrow\mathbb{R}$ with a fixed image $I_{F}$ : $\Omega_{F}\rightarrow\mathbb{R}$. Both images are $D$-dimensional, with each defined over its own spatial domain $\Omega_{F},\Omega_{M}\in\mathbb{R}^{D}$. The application of the transformation to the moving image $I_{M}$ is expressed through function composition: $I_{M}\circ T_{\theta}$. The goal is to find the optimal transformation by minimizing a cost function $\mathcal{C}$ (2), which depends on a similarity metric $\mathcal{S}$ and a $\gamma$-weighted regularization term $\mathcal{P}$. As the central component of the registration process, $\mathcal{S}$ quantifies the alignment quality between the transformed moving image and the fixed image. Its design is crucial, as it controls the robustness, accuracy, and generalization capability of the registration framework, making it the core focus of this study. To complement this, the regularization term $\mathcal{P}(T_{\theta})$ enforces spatial smoothness by penalizing gradients or curvature (bending energy) of the transformation, ensuring a trade-off between deformation flexibility and anatomical plausibility [27].
@@ III-A5 IMPACT : Jacobian mode
For each iteration, $N$ patches of dimension $D$ are extracted uniformly at random. Each patch has a size $P=(P_{1},P_{2},\dots,P_{D})$, where $P_{i}\in\mathbb{N}^{D}$, corresponding to the receptive field at the deeper layer, and is sampled from both the fixed and moving images. Based on these patches, the similarity measure is defined as follows:
@@ III-A6 IMPACT : Static mode
where $\phi_{IF_{l}}:\Omega_{F}\rightarrow\mathbb{R^{C}}$ and $\phi_{IM_{l}}:\Omega_{M}\rightarrow\mathbb{R^{C}}$ represents respectively, from the fixed and moving images, the $C$ features maps extracted at the layer $l$ and $\gamma_{l}$ represent the weight assigned to that layer. The gradient expression of the similarity measure (Eq.8) is simplified as there is no need to compute the Jacobian of $\phi$.
@@ III-B Integration of IMPACT in Elastix Framework
Please refer to Appendix -B for a complete list of metric parameters.
@@ III-C Semantic Similarity Metric Using Pretrained Models for Deep Learning-Based Registration
IMPACT was integrated into the VoxelMorph framework [26] to optimize the learning of a deformable transformation model. VoxelMorph is an unsupervised UNet-based architecture designed to compute a deformation vector field (DVF) for aligning two input images. Its encoder-decoder with skip connections structure allows the encoder to extract multiscale features, while the decoder predicts the DVF. The semantic loss function used in this integration provided two distinct types of features: encoder features, offering general and abstract image representations, and decoder features, optimized for segmentation relevant tasks. Regularization of the DVF is achieved through a diffusion regularizer applied to the spatial gradients, ensuring smooth and diffeomorphic transformations that preserve anatomical topology and continuity. Compared to the Elastix integration, VoxelMorph offers greater flexibility, enabling access to both encoder and decoder features in Jacobian mode while maintaining computational efficiency. This adaptability facilitates efficient exploration of feature representations. This approach can incorporate embeddings from the same broad set of pretrained models (TotalSegmentator, SAM/MedSAM, STU-Net) without modifying VoxelMorph main architecture.
@@ IV-A Clinical tasks
    Thorax FBCT/CBCT registration (ThoraxCBCT Learn2Reg 2023 MICCAI Challenge [58, 12]): This task focuses on thoracic image registration between pre-therapeutic fan beam CT (FBCT) and interventional CBCT scans [58]. The dataset includes three patients, each with longitudinal imaging acquired at different stages of image-guided radiation therapy. For each patient, one FBCT scan is obtained before treatment during maximum inspiration, and two CBCT scans are acquired during maximum expiration. The CBCT images contain substantial artifacts such as scatter, noise, and truncation, which reflect real-world clinical conditions and increase the complexity of the registration task. Evaluation is based on ground truth segmentation masks for lung lobes, tumors, and several organs at risk, including the heart, spinal cord, and esophagus, these annotations are not publicly accessible.
@@ IV-B Common configuration
The IMPACT metric was integrated as a custom similarity measure and configured with task-specific hyperparameters. Depending on the task, features were extracted from different layers of various pretrained models, with the specific combinations detailed in Appendix -A, and compared using different feature distances. In the Elastix framework, two modes of operation were used: Jacobian mode, which enables backpropagation through the feature extractor, and Static mode, which relies on precomputed features and is better suited for extracting deeper semantic representations. The detailed configurations applied to each task are summarized in Table I. All input images were intensity normalized and resampled into a canonical orientation to ensure compatibility with the pretrained models used for semantic feature extraction, following the specifications detailed in Appendix -A.
@@ IV-C Elastix setup
All Elastix-based experiments shared a common configuration for deformable registration. A 3D B-spline transformation model was optimized using the ASGD optimizer [47]. The similarity measure was computed at each iteration over a set of 2000 points randomly sampled in physical coordinates. The use of 2000 sampled points is sufficient to provide a reliable estimate of image similarity and to ensure stable convergence. This was combined with a four-level multi-resolution strategy, where image resolution was progressively refined through resampling at spacings of 6 mm, 3 mm, 1.5 mm, and 1 mm. Simultaneously, the B-spline control point spacing was halved at each resolution level, reaching a final spacing of 8 mm at the finest resolution. This configuration was chosen as a robust default across tasks and is further justified by the analysis in Appendix -C.
@@ IV-E Evaluation Protocol
For task 1, the efficiency of IMPACT was demonstrated by comparing its performance in terms of Target Registration Error (TRE), Dice Similarity Coefficient (DSC) and the 95th percentile Hausdorff distance (HD95) against various conventional similarity measures, including MSE, NCC, NMI and MIND. For a fair comparison, an extraction model for the MIND features has been developed, enabling it to be used in algorithmic registration with Elastix. Four MIND configurations were tested with a radius of 1 and 2, as well as a dilation of 1 and 2. Only the best-performing configuration is presented. Moreover, the convergence speed based on the number of iterations was evaluated, exploring several values: [50, 100, 150, 200, 500, 1000, 1500, 2000]. Task 2 and 3 were evaluated primarily using challenge-specific metrics, including the DSC, HD95, and TRE, which was used exclusively for Task 2. For Task 4, registration accuracy was assessed using only the DSC.
@@ V-B Quantitative evaluation and convergence analysis
TABLE VI: Ranking results of the 32 participants in Task 2 (LungCT) of the Learn2Reg 2021 Challenge. The table presents results based on TRE. Elastix with IMPACT ranked 3th in this leaderboard.
@@
1st (Challenge Organizers) (Fraunhofer MEVIS) 1.8356 $\pm$ 0.5685
2nd (Challenge Organizers) (corrField) 1.8497 $\pm$ 0.7033
3th Our team (BreizhReg) (Elastix) 1.8986 $\pm$ 0.7207
11th (Challenge Organizers) SLIC-Reg++) 2.6359 $\pm$ 1.0163
21th (Challenge Organizers) (Initial Displacements) 14.6407 $\pm$ 6.0765
We further evaluated Elastix with IMPACT on the LungCT task of the Learn2Reg 2021 Challenge, which focuses on aligning lung CT scans using anatomical landmarks as evaluation targets. As reported in Table VI, our method (team “BreizhReg”) achieved a 3th place ranking out of 32 participants, with a mean TRE of 1.90 mm.
@@ VI-A Impact of the Pretrained Model
Several pretrained models were evaluated to assess their influence on the quality and relevance of extracted features for the registration tasks. The models included ConvNeXt, SAM2.1, MedSAM2, and the TotalSegmentator variants M258, M291, and M730 (see Appendix -A for details on training data and target tasks).
@@ VII Discussion
In contrast, both the architectural design (CNN versus transformer) and the domain on which the models were pretrained had a comparatively limited influence. Despite being trained on natural images, the transformer-based SAM2.1 achieved comparable performance to the CNN-based TotalSegmentator. which was trained on medical images. This finding is significant as it demonstrates that sufficiently expressive architectures can learn semantic representations that are general enough to transfer across domains. Furthermore, ablation studies comparing SAM2.1 and MedSAM fine-tuned on medical images, showed identical performance, indicating that domain adaptation had little impact on feature quality. This underscores that the semantic richness introduced by segmentation-based pretraining plays a more decisive role than the pretraining domain itself. These results suggest that models like SAM2.1 can act as truly general-purpose feature extractors, even when applied beyond their original domain.
@@ VIII Conclusion
Moreover, based on our conclusions, we propose the pretrained model M730 with 2 layers and a L2 features distance measure as the recommended generic feature extractor for IMPACT, balancing robustness and efficiency across different registration tasks. However, if the user wishes to further optimize performance for a specific task, the feature extractors in IMPACT can be replaced with a custom model trained on their own data. This allows for more task-specific features while maintaining the versatility of IMPACT, ensuring seamless integration with frameworks like Elastix and VoxelMorph. This adaptability reinforces IMPACT’s role as a scalable and efficient tool, suitable for both research and clinical applications.

### v3

@@ II-A Image Registration: An Optimization Problem
In unsupervised intensity-based image registration, the objective is to estimate a spatial transformation $T_{\theta}$: $\Omega_{F}\rightarrow\Omega_{M}$ that aligns a moving image $I_{M}$ : $\Omega_{M}\rightarrow\mathbb{R}$ with a fixed image $I_{F}$ : $\Omega_{F}\rightarrow\mathbb{R}$. Both images are $D$-dimensional, defined over their respective spatial domains, $\Omega_{F}\subset\mathbb{R}^{D}and\Omega_{M}\subset\mathbb{R}^{D}$. The application of the transformation to the moving image $I_{M}$ is expressed through function composition: $I_{M}\circ T_{\theta}$. The goal is to find the optimal transformation by minimizing a cost function $\mathcal{C}$ (2), which depends on a similarity metric $\mathcal{S}$ and a $\gamma$-weighted regularization term $\mathcal{P}$. As the central component of the registration process, $\mathcal{S}$ quantifies the alignment quality between the transformed moving image and the fixed image. Its design is crucial, as it controls the robustness, accuracy, and generalization capability of the registration framework, making it the core focus of this study. To complement this, the regularization term $\mathcal{P}(T_{\theta})$ enforces spatial smoothness by penalizing gradients or curvature (bending energy) of the transformation, ensuring a trade-off between deformation flexibility and anatomical plausibility [27].
@@ III-A5 IMPACT : Jacobian mode
For each iteration, $N$ patches of dimension $D$ are extracted uniformly at random. Each patch has a size $P=(P_{1},P_{2},\dots,P_{D})$, with $P\in\mathbb{N}^{D}$, corresponding to the receptive field at a deeper layer. For each patch extracted from the fixed image, a corresponding patch is sampled from the moving image using the current spatial transformation. Based on these patches, the similarity measure is defined as follows:
@@ III-A6 IMPACT : Static mode
where $\phi_{IF_{l}}:\Omega_{F}\rightarrow\mathbb{R^{C}}$ and $\phi_{IM_{l}}:\Omega_{M}\rightarrow\mathbb{R^{C}}$ represent respectively, from the fixed and moving images, the $C$ features maps extracted at the layer $l$ and $\gamma_{l}$ represent the weight assigned to that layer. The gradient expression of the similarity measure (Eq.8) is simplified as there is no need to compute the Jacobian of $\phi$.
@@ III-B Integration of IMPACT in Elastix Framework
Please refer to Appendix B for a complete list of metric parameters.
@@ III-C Semantic Similarity Metric Using Pretrained Models for Deep Learning-Based Registration
IMPACT was integrated into the VoxelMorph framework [26] to optimize the learning of a deformable transformation model. VoxelMorph is an unsupervised UNet-based architecture designed to compute a deformation vector field (DVF) for aligning two input images. Its encoder-decoder with skip connections structure allows the encoder to extract multiscale features, while the decoder predicts the DVF. The semantic loss function used in this integration provided two distinct types of features: encoder features, offering general and abstract image representations, and decoder features, optimized for segmentation relevant tasks. Regularization of the DVF is achieved through a diffusion regularizer applied to the spatial gradients, ensuring smooth and diffeomorphic transformations that preserve anatomical topology and continuity. Compared to the Elastix integration, VoxelMorph offers greater flexibility, enabling access to both encoder and decoder features in Jacobian mode while maintaining computational efficiency. This adaptability facilitates efficient exploration of feature representations. This approach can incorporate embeddings from the same broad set of pretrained models (TotalSegmentator, SAM/MedSAM, STU-Net) without modifying VoxelMorph’s main architecture.
@@ IV-A Clinical tasks
    Thorax FBCT/CBCT registration (ThoraxCBCT Learn2Reg 2023 MICCAI Challenge [58, 12]): This task focuses on thoracic image registration between pre-therapeutic fan beam CT (FBCT) and interventional CBCT scans [58]. The dataset includes three patients, each with longitudinal imaging acquired at different stages of image-guided radiation therapy. For each patient, one FBCT scan is obtained before treatment during maximum inspiration, and two CBCT scans are acquired during maximum expiration. The CBCT images contain substantial artifacts such as scatter, noise, and truncation, which reflect real-world clinical conditions and increase the complexity of the registration task. Evaluation is based on ground truth segmentation masks for lung lobes, tumors, and several organs at risk, including the heart, spinal cord, and esophagus. These annotations are not publicly accessible.
@@ IV-B Common configuration
The IMPACT metric was integrated as a custom similarity measure and configured with task-specific hyperparameters. Depending on the task, features were extracted from different layers of various pretrained models, with the specific combinations detailed in Appendix A, and compared using different feature distances. In the Elastix framework, two modes of operation were used: Jacobian mode, which enables backpropagation through the feature extractor, and Static mode, which relies on precomputed features and is better suited for extracting deeper semantic representations. The detailed configurations applied to each task are summarized in Table I. All input images were intensity normalized and resampled into a canonical orientation to ensure compatibility with the pretrained models used for semantic feature extraction, following the specifications detailed in Appendix A.
@@ IV-C Elastix setup
All Elastix-based experiments shared a common configuration for deformable registration. A 3D B-spline transformation model was optimized using the ASGD optimizer [47]. The similarity measure was computed at each iteration over a set of 2000 points randomly sampled in physical coordinates. The use of 2000 sampled points is sufficient to provide a reliable estimate of image similarity and to ensure stable convergence. This was combined with a four-level multi-resolution strategy, where image resolution was progressively refined through resampling at spacings of 6 mm, 3 mm, 1.5 mm, and 1 mm. Simultaneously, the B-spline control point spacing was halved at each resolution level, reaching a final spacing of 8 mm at the finest resolution. This configuration was chosen as a robust default across tasks and is further justified by the analysis in Appendix C.
@@ IV-E Evaluation Protocol
For task 1, the efficiency of IMPACT was demonstrated by comparing its performance in terms of Target Registration Error (TRE), Dice Similarity Coefficient (DSC) and the 95th percentile Hausdorff distance (HD95) against various conventional similarity measures, including MSE, NCC, NMI and MIND. For a fair comparison, an extraction model for the MIND features has been developed, enabling it to be used in algorithmic registration with Elastix. Four MIND configurations were tested with a radius of 1 and 2, as well as a dilation of 1 and 2. Only the best-performing configuration is presented. Moreover, the convergence speed based on the number of iterations was evaluated, exploring several values: [50, 100, 150, 200, 500, 1000, 1500, 2000]. Tasks 2 and 3 were evaluated primarily using challenge-specific metrics, including the DSC, HD95, and TRE, which was used exclusively for Task 2. For Task 4, registration accuracy was assessed using only the DSC.
@@ V-B Quantitative evaluation and convergence analysis
TABLE VI: Ranking results of the 32 participants in Task 2 (LungCT) of the Learn2Reg 2021 Challenge. The table presents results based on TRE. Elastix with IMPACT ranked 3rd in this leaderboard.
@@
1st Challenge Organizers (Fraunhofer MEVIS) 1.8356 $\pm$ 0.5685
2nd Challenge Organizers (corrField) 1.8497 $\pm$ 0.7033
3rd Our team (BreizhReg) (Elastix) 1.8986 $\pm$ 0.7207
11th Challenge Organizers SLIC-Reg++) 2.6359 $\pm$ 1.0163
21th Challenge Organizers (Initial Displacements) 14.6407 $\pm$ 6.0765
We further evaluated Elastix with IMPACT on the LungCT task of the Learn2Reg 2021 Challenge, which focuses on aligning lung CT scans using anatomical landmarks as evaluation targets. As reported in Table VI, our method (team “BreizhReg”) achieved a 3rd place ranking out of 32 participants, with a mean TRE of 1.90 mm.
@@ VI-A Impact of the Pretrained Model
Several pretrained models were evaluated to assess their influence on the quality and relevance of extracted features for the registration tasks. The models included ConvNeXt, SAM2.1, MedSAM2, and the TotalSegmentator variants M258, M291, and M730 (see Appendix A for details on training data and target tasks).
@@ VII Discussion
In contrast, both the architectural design (CNN versus transformer) and the domain on which the models were pretrained had a comparatively limited influence. Despite being trained on natural images, the transformer-based SAM2.1 achieved comparable performance to the CNN-based TotalSegmentator, which was trained on medical images. This finding is significant as it demonstrates that sufficiently expressive architectures can learn semantic representations that are general enough to transfer across domains. Furthermore, ablation studies comparing SAM2.1 and MedSAM fine-tuned on medical images showed identical performance, indicating that domain adaptation had little impact on feature quality. This underscores that the semantic richness introduced by segmentation-based pretraining plays a more decisive role than the pretraining domain itself. These results suggest that models like SAM2.1 can act as truly general-purpose feature extractors, even when applied beyond their original domain.
@@ VIII Conclusion
Moreover, based on our conclusions, we propose the pretrained model M730 with 2 layers and L2 features distance measure as the recommended generic feature extractor for IMPACT, balancing robustness and efficiency across different registration tasks. However, if the user wishes to further optimize performance for a specific task, the feature extractors in IMPACT can be replaced with a custom model trained on their own data. This allows for more task-specific features while maintaining the versatility of IMPACT, ensuring seamless integration with frameworks like Elastix and VoxelMorph. This adaptability reinforces IMPACT’s role as a scalable and efficient tool, suitable for both research and clinical applications.
@@ Acknowledgment
Appendix A Hyperparameters of the IMPACT Similarity Metric
Table XI lists the hyperparameters available for configuring the IMPACT metric. These settings govern the sampling strategy, multi-resolution scheme, feature extraction, and distance computation. Unless stated otherwise, the same parameters were used for both fixed and moving images.
TABLE XI: Hyperparameters for the IMPACT metric used in Elastix-based registration. Parameters cover spatial sampling, feature extraction, and similarity computation.
Parameter Description Default value
MaximumNumberOfIterations Number of iterations for the optimization process at each resolution. 500
NumberOfSpatialSamples Number of spatial samples used during optimization. 2000
NumberOfResolutions Number of multi-resolution levels used in the registration process. 3
FinalGridSpacingInPhysicalUnits Grid spacing for the final resolution in physical units. 8
ModelsPath Path to the pretrained models used for feature extraction. ”Path”
Dimension Dimensionality of the input images used by the model (2D or 3D). 3
NumberOfChannels Number of input channels in the model (like grayscale = 1, RGB = 3). 1
PatchSize Size of the patches extracted for processing. 5*5*5
VoxelSize Resampled voxel size for input images. 1.5*1.5*1.5
LayersMask Binary mask indicating which feature extractor layers are selected. 1
SubsetFeatures Number of selected feature channels. 32
LayersWeight Weight assigned to the selected layers during feature extraction. 1
Mode Mode selection for feature extraction (Static, Jacobian). Jacobian
GPU GPU configuration (CPU mode (-1) or specific GPU device selection). -1
FeaturesMapUpdateInterval Frequency of feature map updates during optimization. -1
PCA Number of principal components retained for the PCA. 0
Loss Choose loss comparison for each layer (L1, L2, NCC, Cosine, L1Cosine). L2
Choice of Multi-Resolution Strategy with Extracted Features
TABLE XII: Comparison of multi-resolution strategies and their impact on registration accuracy (measured with TRE) on Task 1 after 500 iterations.
Multi-resolution strategy TRE (mm)
Jacobian Static
Gaussian smoothing + downsampling (standard multi-resolution) 1.20 1.22
Downsampling only (no smoothing) 1.24 1.24
Gaussian smoothing only (no downsampling) 1.43 1.46
No image pyramid (full-resolution only) 1.94 3.27
We conducted an ablation study on Task 1 (Thorax CT/CBCT) to assess the impact of various multi-resolution strategies on registration accuracy using deep features for similarity computation. Table XII shows the median TRE after 500 iterations with the second-layer model M258, employing L2 distance for feature similarity computation, across four configurations: downsampling only, Gaussian smoothing only, a combination of both (standard multi-resolution), and no pyramid (full resolution only).
The results show that combining Gaussian smoothing and downsampling yields the most accurate performance, especially in Jacobian mode. Additionally, downsampling alone, without smoothing, still produces competitive results, suggesting that spatial rescaling is sufficient on its own. In contrast, operating at full resolution (without the pyramid) results in significantly poorer performance. These findings highlight the advantage of using a four-level multi-resolution strategy, which combines both smoothing and downsampling, as implemented in all Elastix-based experiments.